The textcomp package for using Text Companion fonts ^{∗}

Sebastian Rahtz s.rahtz@elsevier.co.uk 1995/12/11

∗This file has version number 1.4, last revised 1995/12/11.

June 23, 1996

1 Introduction

This style is no more than a formalisation of the work by Jörg Knappen on the text companion fonts.

2 Symbols

1 ⟨∗sty⟩

 2 \DeclareFontEncoding{TS1}{}{}
 3 \DeclareFontSubstitution{TS1}{cmr}{m}{n}

basesingle straight quote

 4 \DeclareTextSymbol{\textquotesinglbase}{TS1}{'015}
 5 \DeclareTextSymbolDefault{\textquotesinglbase}{TS1}

base double straight quotes

 6 \DeclareTextSymbol{\textquotedblbase}{TS1}{'022}
 7 \DeclareTextSymbolDefault{\textquotedblbase}{TS1}

twelve u dash

 8 \DeclareTextSymbol{\texttwelveudash}{TS1}{'025}
 9 \DeclareTextSymbolDefault{\texttwelveudash}{TS1}

threequarters emdash

10 \DeclareTextSymbol{\textthreequartersemdash}{TS1}{'026}
11 \DeclareTextSymbolDefault{\textthreequartersemdash}{TS1}

left pointing arrow

12 \DeclareTextSymbol{\textleftpointingarrow}{TS1}{'030}
13 \DeclareTextSymbolDefault{\textleftpointingarrow}{TS1}

right pointing arrow

14 \DeclareTextSymbol{\textrightpointingarrow}{TS1}{'031}
15 \DeclareTextSymbolDefault{\textrightpointingarrow}{TS1}

blank symbol

16 \DeclareTextSymbol{\textblank}{TS1}{'040}
17 \DeclareTextSymbolDefault{\textblank}{TS1}

dollar sign

18 \DeclareTextSymbol{\textdollar}{TS1}{'044}
19 \DeclareTextSymbolDefault{\textdollar}{TS1}

straight quote

20 \DeclareTextSymbol{\textquotesingle}{TS1}{'047}
21 \DeclareTextSymbolDefault{\textquotesingle}{TS1}

centered star

22 \DeclareTextSymbol{\textcenteredstar}{TS1}{'052}
23 \DeclareTextSymbolDefault{\textcenteredstar}{TS1}

fraction

24 \DeclareTextSymbol{\fraction}{TS1}{'057}
25 \DeclareTextSymbolDefault{\fraction}{TS1}

oldstyle digit 0

26 \DeclareTextSymbol{\textzerooldstyle}{TS1}{'060}
27 \DeclareTextSymbolDefault{\textzerooldstyle}{TS1}

oldstyle digit 1

28 \DeclareTextSymbol{\textoneoldstyle}{TS1}{'061}
29 \DeclareTextSymbolDefault{\textoneoldstyle}{TS1}

oldstyle digit 2

30 \DeclareTextSymbol{\texttwooldstyle}{TS1}{'062}
31 \DeclareTextSymbolDefault{\texttwooldstyle}{TS1}

oldstyle digit 3

32 \DeclareTextSymbol{\textthreeoldstyle}{TS1}{'063}
33 \DeclareTextSymbolDefault{\textthreeoldstyle}{TS1}

oldstyle digit 4

34 \DeclareTextSymbol{\textfouroldstyle}{TS1}{'064}
35 \DeclareTextSymbolDefault{\textfouroldstyle}{TS1}

oldstyle digit 5

36 \DeclareTextSymbol{\textfiveoldstyle}{TS1}{'065}
37 \DeclareTextSymbolDefault{\textfiveoldstyle}{TS1}

oldstyle digit 6

38 \DeclareTextSymbol{\textsixoldstyle}{TS1}{'066}
39 \DeclareTextSymbolDefault{\textsixoldstyle}{TS1}

oldstyle digit 7

40 \DeclareTextSymbol{\textsevenoldstyle}{TS1}{'067}
41 \DeclareTextSymbolDefault{\textsevenoldstyle}{TS1}

oldstyle digit 8

42 \DeclareTextSymbol{\texteightoldstyle}{TS1}{'070}
43 \DeclareTextSymbolDefault{\texteightoldstyle}{TS1}

oldstyle digit 9

44 \DeclareTextSymbol{\textnineoldstyle}{TS1}{'071}
45 \DeclareTextSymbolDefault{\textnineoldstyle}{TS1}

mho sign

46 \DeclareTextSymbol{\textmho}{TS1}{'115}
47 \DeclareTextSymbolDefault{\textmho}{TS1}

big circle

48 \DeclareTextSymbol{\textbigcircle}{TS1}{'117}
49 \DeclareTextSymbolDefault{\textbigcircle}{TS1}
50 \DeclareTextCommand{\textcircled}{TS1}[1]{{%
51 \ooalign{%
52  \hfil \raise .07ex\hbox #1\hfil \crcr
53  \char'117 %

54 }%
55 }}

ohm sign

56 \DeclareTextSymbol{\textohm}{TS1}{'127}
57 \DeclareTextSymbolDefault{\textohm}{TS1}

arrow up

58 \DeclareTextSymbol{\textarrowup}{TS1}{'136}
59 \DeclareTextSymbolDefault{\textarrowup}{TS1}

arrow down

60 \DeclareTextSymbol{\textarrowdown}{TS1}{'137}
61 \DeclareTextSymbolDefault{\textarrowdown}{TS1}

backtick(ASCIIgrave)

62 \DeclareTextSymbol{\textgrave}{TS1}{'140}
63 \DeclareTextSymbolDefault{\textgrave}{TS1}

born

64 \DeclareTextSymbol{\textborn}{TS1}{'142}
65 \DeclareTextSymbolDefault{\textborn}{TS1}

died

66 \DeclareTextSymbol{\textdied}{TS1}{'144}
67 \DeclareTextSymbolDefault{\textdied}{TS1}

divorced

68 \DeclareTextSymbol{\textdivorced}{TS1}{'143}
69 \DeclareTextSymbolDefault{\textdivorced}{TS1}

leaf

70 \DeclareTextSymbol{\textleaf}{TS1}{'154}
71 \DeclareTextSymbolDefault{\textleaf}{TS1}

married

72 \DeclareTextSymbol{\textmarried}{TS1}{'155}
73 \DeclareTextSymbolDefault{\textmarried}{TS1}

musical note

74 \DeclareTextSymbol{\textmusicalnote}{TS1}{'156}
75 \DeclareTextSymbolDefault{\textmusicalnote}{TS1}

low tilde

76 \DeclareTextSymbol{\textlowtilde}{TS1}{'176}
77 \DeclareTextSymbolDefault{\textlowtilde}{TS1}

short equals

78 \DeclareTextSymbol{\textshortequals}{TS1}{'055}
79 \DeclareTextSymbolDefault{\textshortequals}{TS1}

doubletick (ASCII doubleacute)

80 \DeclareTextSymbol{\textquotedbl}{TS1}{'202}
81 \DeclareTextSymbolDefault{\textquotedbl}{TS1}

doublebacktick

82 \DeclareTextSymbol{\textdoublebacktick}{TS1}{'203}
83 \DeclareTextSymbolDefault{\textdoublebacktick}{TS1}

dagger

84 \DeclareTextSymbol{\textdagger}{TS1}{'204}
85 \DeclareTextSymbolDefault{\textdagger}{TS1}

ddagger

86 \DeclareTextSymbol{\textdaggerdbl}{TS1}{'205}
87 \DeclareTextSymbolDefault{\textdaggerdbl}{TS1}

doublevert

88 \DeclareTextSymbol{\textdoublevert}{TS1}{'206}
89 \DeclareTextSymbolDefault{\textdoublevert}{TS1}

perthousand

90 \DeclareTextSymbol{\textperthousand}{TS1}{'207}
91 \DeclareTextSymbolDefault{\textperthousand}{TS1}

bullet

92 \DeclareTextSymbol{\textbullet}{TS1}{'210}
93 \DeclareTextSymbolDefault{\textbullet}{TS1}

centigrade

94 \DeclareTextSymbol{\textcentigrade}{TS1}{'211}
95 \DeclareTextSymbolDefault{\textcentigrade}{TS1}

dollar oldstyle

96 \DeclareTextSymbol{\textdollaroldstyle}{TS1}{'212}
97 \DeclareTextSymbolDefault{\textdollaroldstyle}{TS1}

cent oldstyle

98 \DeclareTextSymbol{\textcentoldstyle}{TS1}{'213}
99 \DeclareTextSymbolDefault{\textcentoldstyle}{TS1}

florin

100 \DeclareTextSymbol{\textflorin}{TS1}{'214}
101 \DeclareTextSymbolDefault{\textflorin}{TS1}

colon

102 \DeclareTextSymbol{\textcolonmonetary}{TS1}{'215}
103 \DeclareTextSymbolDefault{\textcolonmonetary}{TS1}

won

104 \DeclareTextSymbol{\textwon}{TS1}{'216}
105 \DeclareTextSymbolDefault{\textwon}{TS1}

naira

106 \DeclareTextSymbol{\textnaira}{TS1}{'217}
107 \DeclareTextSymbolDefault{\textnaira}{TS1}

guarani

108 \DeclareTextSymbol{\textguarani}{TS1}{'220}
109 \DeclareTextSymbolDefault{\textguarani}{TS1}

peso

110 \DeclareTextSymbol{\textpeso}{TS1}{'221}
111 \DeclareTextSymbolDefault{\textpeso}{TS1}

lira

112 \DeclareTextSymbol{\textlira}{TS1}{'222}
113 \DeclareTextSymbolDefault{\textlira}{TS1}

recipe

114 \DeclareTextSymbol{\textrecipe}{TS1}{'223}
115 \DeclareTextSymbolDefault{\textrecipe}{TS1}

interrobang

116 \DeclareTextSymbol{\textinterrobang}{TS1}{'224}
117 \DeclareTextSymbolDefault{\textinterrobang}{TS1}

gnaborretni

118 \DeclareTextSymbol{\textgnaborretni}{TS1}{'225}
119 \DeclareTextSymbolDefault{\textgnaborretni}{TS1}

vietnamese currency, dong

120 \DeclareTextSymbol{\textdong}{TS1}{'226}
121 \DeclareTextSymbolDefault{\textdong}{TS1}

trademark

122 \DeclareTextSymbol{\texttrademark}{TS1}{'227}
123 \DeclareTextSymbolDefault{\texttrademark}{TS1}

per ten thousand

124 \DeclareTextSymbol{\textpertenthousand}{TS1}{'230}
125 \DeclareTextSymbolDefault{\textpertenthousand}{TS1}

pilcrow

126 \DeclareTextSymbol{\textpilcrow}{TS1}{'231}
127 \DeclareTextSymbolDefault{\textpilcrow}{TS1}

baht (thailand currency)

128 \DeclareTextSymbol{\textbaht}{TS1}{'232}
129 \DeclareTextSymbolDefault{\textbaht}{TS1}

cent

130 \DeclareTextSymbol{\textcent}{TS1}{'242}
131 \DeclareTextSymbolDefault{\textcent}{TS1}

sterling

132 \DeclareTextSymbol{\textsterling}{TS1}{'243}
133 \DeclareTextSymbolDefault{\textsterling}{TS1}

currency sign

134 \DeclareTextSymbol{\textcurrency}{TS1}{'244}
135 \DeclareTextSymbolDefault{\textcurrency}{TS1}

yen

136 \DeclareTextSymbol{\textyen}{TS1}{'245}
137 \DeclareTextSymbolDefault{\textyen}{TS1}

broken vertical bar

138 \DeclareTextSymbol{\textbrokenbar}{TS1}{'246}
139 \DeclareTextSymbolDefault{\textbrokenbar}{TS1}

section sign

140 \DeclareTextSymbol{\textsection}{TS1}{'247}
141 \DeclareTextSymbolDefault{\textsection}{TS1}

high dieresis

142 \DeclareTextSymbol{\texthighdieresis}{TS1}{'250}
143 \DeclareTextSymbolDefault{\texthighdieresis}{TS1}

copyright

144 \DeclareTextSymbol{\textcopyright}{TS1}{'251}
145 \DeclareTextSymbolDefault{\textcopyright}{TS1}

feminineordinalindicator

146 \DeclareTextSymbol{\textordfeminine}{TS1}{'252}
147 \DeclareTextSymbolDefault{\textordfeminine}{TS1}

logical not

148 \DeclareTextSymbol{\textlogicalnot}{TS1}{'254}
149 \DeclareTextSymbolDefault{\textlogicalnot}{TS1}

circled R

150 \DeclareTextSymbol{\textregistered}{TS1}{'256}
151 \DeclareTextSymbolDefault{\textregistered}{TS1}

macron

152 \DeclareTextSymbol{\textmacron}{TS1}{'257}
153 \DeclareTextSymbolDefault{\textmacron}{TS1}

degree sign

154 \DeclareTextSymbol{\textdegree}{TS1}{'260}
155 \DeclareTextSymbolDefault{\textdegree}{TS1}

plus-minus sign

156 \DeclareTextSymbol{\textplusminus}{TS1}{'261}
157 \DeclareTextSymbolDefault{\textplusminus}{TS1}

superscript 2

158 \DeclareTextSymbol{\texttwosuperior}{TS1}{'262}
159 \DeclareTextSymbolDefault{\texttwosuperior}{TS1}

superscript 3

160 \DeclareTextSymbol{\textthreesuperior}{TS1}{'263}
161 \DeclareTextSymbolDefault{\textthreesuperior}{TS1}

tick (ASCII-style acute)

162 \DeclareTextSymbol{\textacute}{TS1}{'264}
163 \DeclareTextSymbolDefault{\textacute}{TS1}

micro sign

164 \DeclareTextSymbol{\textmu}{TS1}{'265}
165 \DeclareTextSymbolDefault{\textmu}{TS1}

pilcrow sign

166 \DeclareTextSymbol{\textparagraph}{TS1}{'266}
167 \DeclareTextSymbolDefault{\textparagraph}{TS1}

centered dot

168 \DeclareTextSymbol{\textperiodcentered}{TS1}{'267}
169 \DeclareTextSymbolDefault{\textperiodcentered}{TS1}

superscript 1

170 \DeclareTextSymbol{\textonesuperior}{TS1}{'271}
171 \DeclareTextSymbolDefault{\textonesuperior}{TS1}

masculine ordinal indicator

172 \DeclareTextSymbol{\textordmasculine}{TS1}{'272}
173 \DeclareTextSymbolDefault{\textordmasculine}{TS1}

fraction one quarter

174 \DeclareTextSymbol{\textonequarter}{TS1}{'274}
175 \DeclareTextSymbolDefault{\textonequarter}{TS1}

fraction one half

176 \DeclareTextSymbol{\textonehalf}{TS1}{'275}
177 \DeclareTextSymbolDefault{\textonehalf}{TS1}

fraction three quarters

178 \DeclareTextSymbol{\textthreequarters}{TS1}{'276}
179 \DeclareTextSymbolDefault{\textthreequarters}{TS1}

multiplication sign (times)

180 \DeclareTextSymbol{\textmultiply}{TS1}{'326}
181 \DeclareTextSymbolDefault{\textmultiply}{TS1}

division sign

182 \DeclareTextSymbol{\textdivide}{TS1}{'366}
183 \DeclareTextSymbolDefault{\textdivide}{TS1}

3 Accents

184 \DeclareTextAccent{\t}{TS1}{'032}

3.1 Acute accent for capitals

185 \DeclareTextAccentDefault{\capitalacute}{TS1}
186 \DeclareTextAccent{\capitalacute}{TS1}{'001}

187 %

188 \DeclareTextAccentDefault{\capitalcircumflex}{TS1}
189 \DeclareTextAccent{\circumflex}{TS1}{'002}

190 %

191 \DeclareTextAccentDefault{\capitaltilde}{TS1}
192 \DeclareTextAccent{\capitaltilde}{TS1}{'003}

193 %

194 \DeclareTextAccentDefault{\capitaldieresis}{TS1}
195 \DeclareTextAccent{\capitaldieresis}{TS1}{'004}

196 %

197 \DeclareTextAccentDefault{\capitalhungarumlaut}{TS1}
198 \DeclareTextAccent{\hungarumlaut}{TS1}{'005}

199 %

200 \DeclareTextAccentDefault{\capitalring}{TS1}
201 \DeclareTextAccent{\capitalring}{TS1}{'006}

202 %

203 \DeclareTextAccentDefault{\capitalcaron}{TS1}
204 \DeclareTextAccent{\capitalcaron}{TS1}{'007}

205 %

206 \DeclareTextAccentDefault{\capitalbreve}{TS1}
207 \DeclareTextAccent{\capitalbreve}{TS1}{'0010}

208 %

209 \DeclareTextAccentDefault{\capitalmacron}{TS1}
210 \DeclareTextAccent{\capitalmacron}{TS1}{'0011}

211 %

212 \DeclareTextAccentDefault{\capitaldotaccent}{TS1}

213 \DeclareTextAccent{\capitaldotaccent}{TS1}{'0012}

214 %

215 \DeclareTextAccentDefault{\capitalcedilla}{TS1}
216 \DeclareTextAccent{\capitalcedilla}{TS1}{'0013}

217 %

218 \DeclareTextAccentDefault{\capitalogonek}{TS1}
219 \DeclareTextAccent{\capitalogonek}{TS1}{'0014}

These command can be used by the end user either directly or through definitions of the type

  \DeclareTextCompositeCommand{\'}{T1}{X}{\capitalacute X}

None of the latter definitions are provided by default, since they are probably rarely used.

220 ⟨/sty⟩ 221 ⟨∗tex⟩

222 \catcode`\_=11
223 \catcode`\ =9
224 \catcode`\^^M=9
225 \catcode`\@=11
226 \catcode`\_=11
227 \catcode`\~=10
228 \def\textcomp_encoding#1#2#3{
229 \gdef\latex_encoding{#2}
230 \if_expert_
231  \gdef\font_encoding{#3}
232 \else
233  \gdef\font_encoding{#1}
234 \fi
235 \gdef\expert_font{}
236 \if_expert_
237  \if_file_exists{
238  \font_family\font_weight\raw_shape\raw_variant
239  \expert_encoding\font_width.mtx
240  }\then
241  \gdef\expert_font{
242  \font_family\font_weight\raw_variant\raw_shape
243  \expert_encoding\font_width,

244 }

245  \fi
246 \fi
247 \installfont{
248  \font_family\font_weight\font_shape\font_variant
249  \font_encoding\font_width

250 }{

251  \font_family\font_weight\raw_shape\raw_variant
252  \raw_encoding\font_width,
253  \expert_font
254  textcomp

255 }{

256  \latex_encoding\encoding_shape

257 }{

258  \latex_encoding

259 }{

260  \latex_family

261 }{

262  \latex_weight\latex_width

263 }{

264  \latex_shape
265 }{}

266 }

267 \def\textcomp_weight#1#2{
268 \gdef\font_weight{#1}%
269 \gdef\latex_weight{#2}%
270 \textcomp_widths

271 }

272

273 \def\textcomp_weights{
274 \textcomp_weight{b}{b}
275 \textcomp_weight{d}{db}
276 \textcomp_weight{i}{el}
277 \textcomp_weight{l}{l}
278 \textcomp_weight{s}{sb}
279 \textcomp_weight{k}{m}
280 \textcomp_weight{m}{m}
281 \textcomp_weight{r}{m}
282 \textcomp_weight{x}{eb}
283 \textcomp_weight{u}{ub}

284 }

285

286 \def\textcomp_widths{
287 \textcomp_width{}{}
288 \textcomp_width{c}{sc}

289 }

290

291 \def\textcomp_shapes{
292 \textcomp_shape{} {} {} {n}
293 \textcomp_shape{o}{o}{} {sl}
294 \textcomp_shape{i}{i}{i}{it}

295 }

296

297 \def\textcomp_encodings{
298 \textcomp_encoding{8c}{TS1}{9c}

299 }

300 \def\textcomp_width#1#2{
301 \gdef\font_width{#1}
302 \gdef\latex_width{#2}
303 \textcomp_shapes

304 }

305

306 % textcomp shape{FONTSHAPE}{RAWSHAPE}{ENCODINGSHAPE}{LATEXSHAPE}

307

308 \def\textcomp_shape#1#2#3#4{
309 \gdef\font_shape{#1}
310 \gdef\raw_shape{#2}
311 \gdef\encoding_shape{#3}
312 \gdef\latex_shape{#4}

313 \csname fake_shape_\raw_shape\endcsname
314 \if_file_exists{
315  \font_family\font_weight\raw_shape
316  \raw_variant\raw_encoding\font_width.mtx
317 }\then
318  \textcomp_encodings
319 \fi

320 }

321

322 \def\textcompfamily#1#2{
323 \edef\temp_command{#1}
324 \expandafter\parse_family\temp_command
325  \empty_command\empty_command\end_parse_family
326 \installfonts
327  \installfamily{TS1}{#1}{#2}
328  \if_file_exists{ \raw_encoding.mtx }\then
329  \installfamily{\raw_encoding}{#1}{#2}
330  \fi
331  \textcomp_weights
332 \endinstallfonts

333 } 334 ⟨/tex⟩ 335 ⟨∗test⟩

336 \documentclass{article}
337 \usepackage{textcomp}
338 \usepackage[T1]{fontenc}
339 \def\rmdefault{cmr}
340 \begin{document}
341 \begin{tabular}{|ll|ll|}
342 \hline
343 fraction & \fraction & textacute & \textacute \\
344 textarrowdown & \textarrowdown & textarrowup & \textarrowup \\
345 textbaht & \textbaht & textbigcircle & \textbigcircle \\
346 textblank & \textblank & textblank  & \textblank  \\
347 textborn & \textborn & textbrokenbar & \textbrokenbar \\
348 textbullet & \textbullet & textcent & \textcent \\
349 textcenteredstar & \textcenteredstar & textcentigrade & \textcentigrade \\
350 textcentoldstyle & \textcentoldstyle & textcolonmonetary & \textcolonmonetary \\
351 textcopyright & \textcopyright & textcurrency & \textcurrency \\
352 textdagger & \textdagger & textdaggerdbl & \textdaggerdbl \\
353 textdegree & \textdegree & textdied & \textdied \\
354 textdivide & \textdivide & textdivorced & \textdivorced \\
355 textdollar & \textdollar & textdollaroldstyle & \textdollaroldstyle \\
356 textdong  & \textdong  & textdoublebacktick & \textdoublebacktick \\
357 textdoublevert & \textdoublevert & texteightoldstyle & \texteightoldstyle \\
358 textfiveoldstyle & \textfiveoldstyle & textflorin & \textflorin \\
359 textfouroldstyle & \textfouroldstyle & textgnaborretni & \textgnaborretni \\
360 textgrave & \textgrave & textguarani & \textguarani \\
361 texthighdieresis & \texthighdieresis & textinterrobang & \textinterrobang \\
362 textleaf & \textleaf & textleftpointingarrow & \textleftpointingarrow \\
363 textlira & \textlira & textlogicalnot & \textlogicalnot \\
364 textlowtilde & \textlowtilde & textmacron & \textmacron \\
365 textmarried & \textmarried & textmho & \textmho \\
366 textmu & \textmu & textmultiply & \textmultiply \\

367 textmusicalnote & \textmusicalnote & textnaira & \textnaira \\
368 textnineoldstyle & \textnineoldstyle & textohm & \textohm \\
369 textonehalf & \textonehalf & textoneoldstyle & \textoneoldstyle \\
370 textonequarter & \textonequarter & textonesuperior & \textonesuperior \\
371 textordfeminine & \textordfeminine & textordmasculine & \textordmasculine \\
372 textparagraph & \textparagraph & textperiodcentered & \textperiodcentered \\
373 textpertenthousand & \textpertenthousand & textperthousand & \textperthousand \\
374 textpeso & \textpeso & textpilcrow & \textpilcrow \\
375 textplusminus & \textplusminus & textquotedbl & \textquotedbl \\
376 textquotedblbase & \textquotedblbase & textquotesinglbase & \textquotesinglbase \\
377 textquotesingle & \textquotesingle & textrecipe & \textrecipe \\
378 textregistered & \textregistered & textrightpointingarrow & \textrightpointingarrow \\
379 textsection & \textsection & textsevenoldstyle & \textsevenoldstyle \\
380 textshortequals & \textshortequals & textsixoldstyle & \textsixoldstyle \\
381 textsterling & \textsterling & textthreeoldstyle & \textthreeoldstyle \\
382 textthreequarters & \textthreequarters & textthreequartersemdash & \textthreequartersemdash \\
383 textthreesuperior & \textthreesuperior & texttrademark & \texttrademark \\
384 texttwelveudash & \texttwelveudash & texttwooldstyle & \texttwooldstyle \\
385 texttwosuperior & \texttwosuperior & textwon & \textwon \\
386 textyen & \textyen & textzerooldstyle & \textzerooldstyle \\
387 \hline
388 \end{tabular}
389 \end{document}

390 ⟨/test⟩